OpenStudy (anonymous):

can you check if im correct plzz Renee and Zelda are at position A. At the same time, they start moving toward their destination B at time t = 0. Renee's speed is given by the function v A = 4 + 3t, and Zelda's speed is given by the function v B = 5 + 2t, where t is the number of seconds and v is meters/second. What is their total speed at t = 3 seconds?

OpenStudy (anonymous):

is the answer 2 meters/second?

OpenStudy (anonymous):

4+3(3)=13 5+2(3)=11 13-11=2

OpenStudy (anonymous):

nvm i was suppose to add so would it be24? bc it's asking for the total speed